OpenStudy (anonymous):

What is the magnitude and direction of if G(2, -2) and H(7, 6)?

OpenStudy (anonymous):

magnitude: 9.43 units; direction 32.01° magnitude: 6.4 units; direction: 57.99° magnitude: 9.43 units: direction: 57.99° magnitude: 6.4 units: direction: 32.01°

OpenStudy (anonymous):

?????????

OpenStudy (anonymous):

why not use length of a line formula ?

OpenStudy (anonymous):

Whats the formula then

OpenStudy (anonymous):

that is the distance between 2 points ., using that you can find the mag

OpenStudy (anonymous):

OOOOOO the distance formula lol

OpenStudy (anonymous):

yea and you can also ., use the gradient formula :)

OpenStudy (anonymous):

to find the direction you kno (y-y1/x-x1 ) ?

OpenStudy (anonymous):

So the answer is

OpenStudy (jdoe0001):

magnitude of a vector <a , b> => \(\bf || <a, b> || = \sqrt{a^2+b^2}\)

OpenStudy (jdoe0001):

what does -> G(2, -2) and H(7, 6) <- mean? G+H => (2, -2) + (7, 6) ?

OpenStudy (anonymous):

So whats the answer

OpenStudy (jdoe0001):

well, do the magnitude calculations

OpenStudy (jdoe0001):

and to get the direction of a vector => \(\bf cos(\theta)=\frac{u \cdot v}{||u||\times ||v||}\)

OpenStudy (anonymous):

ok i found and easier way sry i confused you with the gradient

OpenStudy (anonymous):